# Effects of interlayer Dzyaloshinskii–Moriya interaction on the shape and dynamics of magnetic twin-skyrmions

### v1

@@ Methods
Atomistic spin model
We describe the magnetic system by a classical atomistic spin model on two layers of a square lattice containing $128\times 128$ sites on top of each other. We split the spin Hamiltonian of our system into three parts,
$H=H_{1}+H_{2}+H_{\text{inter}},$ (7)
$-\mu_{s}\sum_{i}\mbox{$\mathrm{B}$}\cdot\mbox{$\mathrm{S}$}_{i}^{l},$ (8)
Inside each layer we consider the nearest-neighbor interfacial (IF) exchange interaction ($J^{\text{IF}}$), the nearest-neighbor DMI ($\mbox{$\mathrm{D}$}^{\text{IF}}_{ij}$) vector perpendicular to the bonds between nearest neighbors preferring a Néel-type rotation, and the Zeeman interaction with an external magnetic field $\mathrm{B}$. $\mu_{s}$ is the magnetic moment at a site and $\langle i,j\rangle$ is the sum over all sites $i$ and their nearest neighbors $j$. We include two interlayer couplings mediated by a non-magnetic spacer layer separating the two magnetic layers: the IEC $J^{\text{IL}}$ and the IL-DMI $\mbox{$\mathrm{D}$}^{\text{IL}}$. If not specified otherwise, we consider the parameters $J^{\text{IF}}=$6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{D}$}^{\text{IF}}|=$1.5\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, and $\mu_{s}=3\mu_{B}$, where $\mu_{B}$ is the Bohr magneton. These values lead to the stabilization of Néel-type isolated metastable skyrmions in a collinear background inside a single layer for a sufficiently high external magnetic field applied along the out-of-plane $z$ direction, as illustrated in Fig. 1(a).
We simulate the time evolution of the spins by numerically solving the Landau-Lifshitz-Gilbert (LLG) equation [41, 42],
$\dot{\mbox{$\mathrm{S}$}}_{i}^{l}=-\frac{\gamma}{1+\alpha^{2}}\mbox{$\mathrm{S}$}_{i}^{l}\times\left(\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}+\alpha\mbox{$\mathrm{S}$}^{l}_{i}\times\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}\right),$ (10)
where $\alpha$ is the Gilbert damping and $\gamma$ is the absolute value of the gyromagnetic ratio. The effective field $\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}$ is given by the derivative of the Hamiltonian with respect to $\mbox{$\mathrm{S}$}^{l}_{i}$,
$\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}=$ $-\frac{1}{\mu_{s}}\frac{\partial H}{\partial\mbox{$\mathrm{S}$}^{l}_{i}}$
$=$ $\frac{J^{\text{IF}}}{\mu_{s}}\sum_{\langle j\rangle_{i}}\mbox{$\mathrm{S}$}_{j}^{l}+\frac{1}{\mu_{s}}\sum_{\langle j\rangle_{i}}\mbox{$\mathrm{D}$}_{ij}^{\text{IF}}\times\mbox{$\mathrm{S}$}_{j}^{l}+\mbox{$\mathrm{B}$}$
@@
Current-driven motion
To study the dynamics of the structures, we consider the current-perpendicular-to-the-plane (CPP) geometry. This describes the case when the multilayer is placed on a non-magnetic substrate with strong spin–orbit coupling. Besides the substrate being responsible for the interfacial DMI, an electric current flowing inside the substrate gives rise to a spin current flowing perpendicular to the layers, which exerts a spin–orbit torque on the magnetization [23]. This torque is included in the LLG equation as follows [43, 18],
where $\mathrm{P}$ is the direction of the current polarization. It will be assumed that the polarization is perpendicular to both the electric current or electric field direction inside the nonmagnetic substrate and the flowing direction of the spin current perpendicular to the surface [23], $\mbox{$\mathrm{P}$}\parallel\mbox{$\mathrm{j}$}_{\textrm{el}}\crossproduct\hat{\mbox{$\mathrm{z}$}}$. The effect of the current is split into a field-like torque proportional to $\beta_{s}$ and a damping-like torque proportional to $\beta_{d}$. Since the field-like torque acts exactly like an external magnetic field, it distorts the skyrmions but does not set them into motion [21]. Therefore, we simplify the analysis by setting $\beta_{f}=0$ and only investigating the effect of the damping-like term driving the skyrmions.
@@
All the data presented in this work are available from the corresponding author upon reasonable request.
@@
All the codes for calculating the data and visualizing it are available from the corresponding author upon reasonable request.
@@ Acknowledgments
The authors declare no competing interests.
@@
Correspondence and requests for materials should be addressed to Elena Y. Vedmedenko.

### v2

@@ Methods
The effective field $\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}$ in the LLG equation (5) is given in the two layers as
$\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}=$ $\frac{J^{\text{IF}}}{\mu_{s}}\sum_{\langle j\rangle_{i}}\mbox{$\mathrm{S}$}_{j}^{l}+\frac{1}{\mu_{s}}\sum_{\langle j\rangle_{i}}\mbox{$\mathrm{D}$}_{ij}^{\text{IF}}\times\mbox{$\mathrm{S}$}_{j}^{l}+\mbox{$\mathrm{B}$}$
@@
The simulation is performed on a $128\times 128$ square lattice with periodic boundary conditions for each layer. The used integration scheme is the Euler method with an additional normalization step after each iteration. The time step is $\Delta t=$1\text{\times}{10}^{-6}\text{\,}\mathrm{n}\mathrm{s}$\$ for Figs. 1, 4, and 5. Some simulations were repeated with various time steps to confirm the stability of the solution for the selected time step. For the static properties in Figs. 2 and 3, a larger time step $\Delta t=$2\text{\times}{10}^{-6}\text{\,}\mathrm{n}\mathrm{s}$\$ was used, since here the focus is on obtaining a local energy minimum instead of following the time evolution. The simulation was always terminated after a fixed number of iterations. The number of iterations is given in each figure caption. Close to the boundary of the stability region in Fig. 2, we confirmed the stability by increasing the simulation time by a factor of ten, and we did not observe a measurable change in the skyrmion size, with this size staying well below the size of the simulation cell. Outside the stability region, we increased the system size by a factor of two, and observed a further elongation of the skyrmion until it filled up the simulation cell and was stopped by confinement.
@@
The datasets generated or analysed during the current study are not publicly available due to it being not technically feasible or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project, but are available from the corresponding author on reasonable request.
@@
The underlying code for this study is not publicly available but may be made available to qualified researchers on reasonable request from the corresponding author.
@@ Acknowledgments
The authors declare no competing financial or non-financial interests.
@@
Correspondence and requests for materials should be addressed to Levente Rózsa or Elena Y. Vedmedenko.